[figure: 9.7 cm]

A MIXED-SPIN INTERFACE CONTRIBUTION TO GMR-EFFECTS IN LAYERED STRUCTURES

G. Wiatrowski^{a}, K. Warda^{a}, L. Wojtczak^{a}, D.Baldomir^{b}, M. Pereiro^{b} and J.E. Arias^{b} aSolid State Physics Department, University of Lodz, Poland bDepartamento de Fisica Aplicada Universidad Santiago de Compostela, Spain.

We discuss a mixed-spin Ising model of thin ferromagnetic layered structures with interface magnetic moment (S^{A} = 1/2, S^{B} = 1), perpendicular uniaxial anisotropy and interface exchange interactions varied. The temperature dependent magnetization profiles, the phase diagrams of critical temperature T_{C} vs. concentration c_{B} of B-atoms and interface interactions J_{AB} and J_{BB} are presented. In particular, the magnetization profiles across layered structure obtained for different interface parameters lead to various interface potentials. The change of potentials across the ferromagnetic layers was applied for calculation of Giant Magnetoresistance effect (GMR) in the frame of the extended Hood-Falicov theory. The curves of GMR versus thickness of ferromagnetic layer with different surface parameters are presented.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : Grzegorz Wiatrowski

Address for correspondence : Department of Solid State Physics University of Łódź Pomorska 149/153 90–236 Łódź

Email address : wiatr@krysia.uni.lodz.pl